OpenStudy (anonymous):

Simplify 2(–3x3y4)2. A. –18x^9y^16 B. 12x^9y^16 C. 18x^6y^8 D. –12x^6y^8

OpenStudy (anonymous):

4(-3x3y4) and distribute or is that second two squaring the expression

OpenStudy (anonymous):

only squaring the equation

OpenStudy (anonymous):

there is something strange about your answers though. i don't know how they would become exponents of themselves.

OpenStudy (anonymous):

I was given these answers... :S

OpenStudy (anonymous):

2 (-3 x^3 y^4)^2

OpenStudy (anonymous):

that is the original expression so foil out the parenthesis part

OpenStudy (anonymous):

and distribute your two and you should get 18 x^6 y^8

OpenStudy (texaschic101):

I agree....18x^6y^8

OpenStudy (anonymous):

\[2 (-3 x^3 y^4)^2 = 2 ((-3 x^3 y^4)(-3 x^3 y^4))\] \[2((9x^6y^8))= 18x^6y^8\]

OpenStudy (anonymous):

Thanks!